OpenStudy (lolbunny):

Which of the following is not a way to represent the solution of the inequality 5(x + 2) greater than or equal to 7x + 2(x − 1)? A number line with a closed circle on 3 and shading to the right A number line with a closed circle on 3 and shading to the left 3 greater than or equal to x x less than or equal to 3

OpenStudy (anonymous):

A number line with a closed circle on 3 and shading to the left is not a way to represent the solution

OpenStudy (anonymous):

5(x+2) >= 7x + 2(x-1) if we simplify by distributing the 5 and two to their corresponding parenthesis we get 5x + 10 >= 7x + 2x -2 We can now add the 7x and 2x 5x + 10 >= 9x - 2 by subtracting 5x on both sides 10 >= 4x -2 adding 2 to both sides 12 >= 4x divide both sides by 4 3 >= x

OpenStudy (anonymous):

3 >= x can be said 3 is greater than or equal to x and can also be said x is less than or equal to x. So those are not the answers

OpenStudy (anonymous):

Remember that when we have a "equals to" there is a shaded circle and in this case we ave a shaded circle on 3. Just look at 3 >= x

OpenStudy (anonymous):

since 3 is greater than or equal to x the number line will look like <---------o==========> where o is a shaded circle and === is shaded to the right 3 this means that A number line with a closed circle on 3 and shading to the right is there

rebeccaxhawaii (rebeccaxhawaii):

solve for x

rebeccaxhawaii (rebeccaxhawaii):

@tkhunny